Two athletes a and b set off from the same point but in opposite directions at a constant speed around a track of circumference 400m.if they meet after 50 s and the speed of a is 1.5 times the speed of b.calculate the speed of each


1
Expert's answer
2021-08-18T02:24:13-0400

Let the speed of athlete B is "v", then speed of of athlete A is "1.5v". We get

"l=v_At+v_Bt=2.5vt"

Hence,

"v=l/2.5t=400\:\rm m/(2.5*50\: s)=3.2\: m/s"

Finally

"v_B=3.2\: {\rm m/s}\\
v_A=1.5*3.2\: \rm m/s=4.8\:\ m/s"